OpenStudy (anonymous):

Find the product of (2x − 5y)^2. 4x^2 + 20xy + 25y^2 4x^2 − 20xy + 25y^2 4x^2 − 25y^2 4x^2 + 25y^2 @eSpeX

OpenStudy (campbell_st):

this is a perfect square so the expansion will follow the one below \[(a - b)^2 = a^2 - 2ab + b^2\]

OpenStudy (anonymous):

How would I set that up?

OpenStudy (precal):

just expand it out, (using FOIL)

OpenStudy (precal):

but if you want to use the formula a is 2x and b is 5y

OpenStudy (espex):

All you need to do is plug your 'a' and 'b' terms into the template @campbell_st gave you.

OpenStudy (campbell_st):

try this method \[(2x - 5y)^2 = (2x -5y)(2x-5y)\] now this can be written as \[2x(2x - 5y) - 5y(2x - 5y)\] just distribute

OpenStudy (anonymous):

http://www.sketchtoy.com/35402938 Is this correct?

OpenStudy (anonymous):

@radar ?

OpenStudy (whpalmer4):

yes, that's correct.

OpenStudy (anonymous):

Thanks!

OpenStudy (whpalmer4):

it all boils down to multiplying every piece of the first thing with every piece of the second thing, then adding the pieces together.